# The universal multiplicity function: counting halos and voids

### v1

@@ 5 Eulerian mapping
i) Our VSF model with the effective moving barrier, Eq. (16), using the exact numerical multiplicity function (solid blue) from Eqs. (5)–(6) and the moving barrier of Eq. (16) with the MCMC best-fit parameters obtained in Lagrangian space, Eq. (17). The blue dashed line shows the corresponding analytical approximation, Eq. (3). Both are spherically mapped in Eulerian space, Eq. (19).
@@
ii) The SVdW model (green curve) with linear formation threshold $\delta_{\rm v}=-0.623$, mapped in Eulerian space via Eq. (19).
@@
iii) The Jennings et al. (2013) VSF model (hereafter Vdn) with linear formation threshold $\delta_{\rm v}=-0.623$ (pink curve). This model, also known as the volume-conserving model, is a Eulerian mapping of the SVdW, which relies mainly on one assumption: the total volume fraction of voids is conserved from Lagrangian to Eulerian space, at the expense of the conservation of the number of voids. This assumption was introduced to avoid normalization problems in the sharp-$k$ multiplicity function (SVdW) spherically mapped to the Eulerian space.
@@
Figure 5: Eulerian HMF. Top panel: HMF measured from DEMNUni at $z=0$ (black dots) and corresponding Poissonian uncertainty (errorbars); Eulerian spherical mapping of the effective barrier approach implemented in this work (blue), Eqs. (3), (16), fitting the halo threshold from Eq. (17); SMT best-fit model (orange); 4-parameters Tinker et al. (2008) best-fit model. Bottom panel: relative difference between the theoretical models of the upper panel (same color labeling) with respect to HMF measured in the DEMNUni in unit of Poissonian uncertainty. Errorbars and hatched area show ±$\pm 1\sigma$ interval, dashed area shows the ±$\pm 1\%$ uncertainty
@@
Finally, the lower panel of Fig. 4 shows the residuals at $z=0$ of the above Eulerian VSF models with respect to VSF measurements from the DEMNUni (dark colors) and HR-DEMNUni (transparent) data (the errorbars represent the Poissonian uncertainty). It can be observed how the Eulerian mapping of SVdW and the Vdn model is not able to reproduce the VSF from simulations. Concerning our model, it can be noticed that the spherical mapping is accurate within $1\sigma$ at both large and intermediate scales, while it becomes less accurate at lower scales, on which the spherical Eulerian map over-predicts the number of voids by about $\sim 50\%$. The accuracy at larger scales is in agreement with the fact that the anisotropic part of the strain of the displacement field around density extremants decreases as the scale increases (Bardeen et al., 1986; Bond et al., 1996). In particular, the expected value of the ellipticity is proportional to $\sigma(R)/|\delta_{\rm v}|$ (SMT), which makes the spherical approximation accurate on large scales. On smaller scales, other effects, such as non-sphericity and non-locality, may become relevant. From the same arguments, it follows that the accuracy of the spherical Eulerian mapping increases with the redshift.
@@
The DM halo case. Lagrangian halos undergo a full collapse in Eulerian space, which makes the mapping conceptually different from the void case. On the one hand, at each redshift it corresponds a unique linear collapse threshold (see Appendix E for details); on the other hand, the spherical symmetric evolution can hardly be accurate in the strongly nonlinear regime of the collapse. In the literature it is shown that, at the first-order approximation, the non-spherical collapse is reflected in a scale-dependent (SMT) and fuzzy formation threshold (Robertson et al., 2009; Ludlow & Porciani, 2011) in Lagrangian space. In addition, non-linear evolution can break the top-hat Lagrangian to Eulerian mass conservation, $M_{\rm E}(R)=M_{\rm TH}(R)$, as the fluid elements of a Lagrangian spherical overdensity may not all and exclusively collapse in the corresponding Eulerian halo (Ludlow & Porciani, 2011). Finally, above all, there is the problem of estimating which regions in the Universe (or in simulations) are collapsed and eventually virialized, i.e. are DM halos. Different prescriptions result in different halo finders (Knebe et al., 2013). This makes an a priori modeling of halos statistics, such as the HMF, extremely challenging. As a consequence, the halo multiplicity function needs to be fitted according to the different halo finder considered (Tinker et al., 2008SMT). In this perspective, we use the HMF measured at $z=0$ from the DEMNUni simulations to find the corresponding halo formation threshold for Lagrangian halos, as defined in Sec. 2. In particular, we consider the halo catalog corresponding to an overdensity threshold equal to the virial value and obtained by exploiting the spherical overdensity package within the Subfind algorithm (Springel et al., 2001; Dolag et al., 2009) in Gadget-3. According to the spherical mapping approximation, we assume the linear halo formation threshold, $\delta_{\rm h}$, to be scale independent and the Eulerian mass to be the same as the Lagrangian one, $M_{\rm E}(R)=M_{\rm TH}(R)$. As explained in Appendix E, to fit the threshold we use Eq. (17) as a prior, after substituting $\delta_{\rm v}$ with $\delta_{\rm h}$. We find a redshift independent best-fit $\delta_{\rm h}=2.424\pm^{0.076}_{0.071}$. Moreover the corresponding effective barrier parameters in Eq. (16) are found to be redshift independent as well, satisfying the regime in which the HMF can be considered self-similar, confirming its redshift universality (see Appendix E for details). Note that, even if this value seems to be larger than the usual spherical collapse threshold, $\delta_{\rm c}\simeq 1.686$, actually our results confirm previous finding in the literature (Robertson et al., 2009). In fact, the barrier value corresponding to $\delta_{\rm h}=2.424\pm^{0.076}_{0.071}$ is $\alpha=1.209\pm^{0.008}_{0.007}$, only a $\sim 15\%$ different from the  SMT moving barrier $\sqrt{0.7}\delta_{\rm c}\sim 1.4$.
@@
Fig. 5 shows the results. The upper panel shows: i) the sub-HMF measured at $z=0$ from the DEMNUni (black dots), with the corresponding Poissonian uncertainty (error bars); ii) threshold fit $\delta_{\rm h}$ of theoretical HMF from the moving barrier approach using Eqs. (3)–(16)–(17) (blue curve); iii) the SMT best-fit model (orange curve); iv) 4-parameters Tinker et al. (2008) best-fit model. The bottom panel shows the relative differences between the measurements from simulations and the best-fit models of the upper panel in units of the Poissonian uncertainty. The errorbars and the hatched area show the ±$\pm 1\sigma$ interval, the dashed area shows the ±$\pm 1\%$ uncertainty. It can be noticed that, on the scales considered, the model presented in this work, Eqs. (3)–(16)–(17), plus the fitting of the linear formation threshold $\delta_{\rm h}$, allow us to reach a better accuracy than for the SMT best-fit model and similar to the 4-parameters Tinker et al. (2008) best-fit model. Appendix E shows the redshift dependence of all these theoretical HMFs.
@@ 6 The void density profile from the effective barrier
In this Section, we explore the effective barrier approach for statistics beyond the multiplicity function, by considering the void density profile of Lagrangian voids. Note that this is the integral of the void-matter cross-correlation function. Here we focus on the density profile of voids because, beyond the motivations explained at the beginning of Sec. 4, halo collapse destroys the one-to-one map between the Lagrangian and Eulerian profile, which is not the case for voids. Moreover, a theoretical modeling of the Lagrangian density profile of voids would be useful for their analysis in the redshift space (Pisani et al., 2014; Hawken et al., 2017; Hamaus et al., 2016; Cai et al., 2016; Radinović et al., 2023).
@@
In the standard excursion-set approach, each stochastic realization of the $\delta({\bf q},R)$ field, given by Eqs. (2), statistically corresponds to a realization of the integrated density profile with respect to a random Lagrangian position ${\bf q}$. Here, we apply this feature using our effective barrier approach to void formation, in order to verify whether it can reproduce the density profile around minima. As the effective barrier $B$ in Eqs. (16)–(17) does not correspond to the void formation threshold $\delta_{\rm v}$, the corresponding density profile realizations would not describe the physical ones. We avoid this effect by considering two steps. First, at scales larger than the crossing radius, $R_{\times}$, we normalize the field to the actual void formation threshold as
$\delta(R)\rightarrow\delta(R)\,\delta_{\rm v}/B(R)\quad{\rm for}\quad R\geq R_{\times}.$ (20)
@@
This step aims to rescale the amplitude of the field from the effective to the physical one. Second, at scales smaller than the crossing radius, we shift the stochastic realization of the field as
$\delta(R)\rightarrow\delta(R)-B(R_{\times})+\delta_{\rm v}\quad{\rm for}\quad R<R_{\times}.$ (21)
@@
This shifts the stochastic realization $\delta(R)$ by an offset chosen to match the physical void formation threshold at the crossing radius, $\delta(R_{\times})=\delta_{\rm v}$. After crossing, the field behaves as a random walk pinned at the coordinate $(R_{\times},\delta_{\rm v})$ in the $R-\delta$ plane and therefore satisfies the above equation. It follows that the width of the distribution of void profiles is approximated by $\sqrt{s-S(R_{\times})}$ as $s$ increases ($R$ decreases).
@@
Figure 6: Lagrangian density profile, linearly evolved at $z=0$, for voids with linear formation threshold of $\delta_{\rm v}=-0.388$ and void radius between $15-21$Mpc (purple), $33-39$Mpc (red), $58-65$Mpc (green), and $(72-79)$ $h^{-1}$pc (blue). The solid lines show the excursion-set prediction with the effective moving barrier obtained as described in 6; the light hatched area the shortest interval subtending 68% of the distribution; the dashed lines show the mean void density profile measured in simulations ICs; the light shaded area the shortest interval subtending 68% of the distribution.
@@
Fig. 6 shows the comparison between the void density profile measured in simulations ICs and linearly evolved at $z=0$ (dashed lines), with respect to stochastic realizations of $\delta(R)$, obtained with the corresponding effective moving barrier, Eqs. (16)–(17), rescaled via Eqs. (20)–(21) (solid lines). Different colors represent voids with different formation radii, as specified in the caption. The shaded areas show the shortest interval subtending $68\%$ of the distribution of density profiles in simulations ICs; while the hatched areas show the same quantity for stochastic realizations. Interestingly, in both cases, the distributions are well described by a Gaussian PDF for each radius $R$. The correspondence between simulations ICs and stochastic realizations is very good, in particular: i) the void profile distribution obtained via the effective barrier approach closely matches the one from simulations ICs, as can be noted by the shaded and hatched areas in Fig. 6; ii) the mean of the excursion-set profile distribution reproduces the global behavior of the mean profile measured in the simulations ICs, while the accuracy depends on the void scale considered; iii) the change of slope around the crossing radius is accurately reproduced; iv) interestingly, the effective barrier predicts the formation of a non-collapsing ovedensity in which smaller voids are embedded, as observed in simulations ICs.
@@
Our void profile prediction in Lagrangian space is the third main result of this work, and the methodology to model it has been presented here for the first time in the literature. In particular, this modeling uses the same effective barrier of Eqs. (16)–(17) used to derive the theoretical multiplicity function and VSF presented in this work. This modeling would not be representative of simulation ICs measurements unless the effective barrier contained statistical information beyond the multipiplicity function.
@@ 7 Conclusions
In this work, we show that using an effective moving barrier in the excursion-set framework, it is possible to combine the latter with the peak theory in Lagrangian space. In particular: i) we provide in Sec. 2 a precise halo and void formation definition in Lagrangian space, that was missing for voids, which combines excursion-set and peak theory; ii) in Sec. 3, for a general moving barrier in the presence of scale correlations, we derive Eqs. (9), (3), and (3), i.e. a multiplicity function able to recover both halo and void statistics from small to very large scales $S$. In particular, Eq. (3) represents a new and very useful accurate analytical approximation of the general solution (see Fig. 1 for a detailed comparison with the numerical multiplicity function obtained via Monte Carlo realizations); iii) in Sec. 4 we derive, in the void case, the parameters characterizing the SMT-like effective moving barrier, in Eq. (16), as a function of the void formation threshold, $\delta_{\rm v}$, by inserting Eq. (16) in Eq. (6) and Eq. (3) and fitting the VSF obtained via the latter against the Lagrangian VSF measured in the DEMNUni ICs; iv) in Sec. 5 we discuss the Lagrangian to Eulerian mapping, in particular we implement the spherical map in Eulerian space for the Lagrangian VSF and HMF, introduced in Secs. 34. We obtain a good agreement with simulations for intermediate and large voids and we accurately reproduce the HMF by fitting the linear halo formation threshold from simulations; v) in Sec. 6 we show that, in the excursion-set framework with the effective barrier, it is possible to derive the Lagrangian void density profile, using the effective barrier of Eqs. (16)–(17) entering the corresponding multiplicity function, Eq. (6) and Eq. (3), derived in this work. See Fig. 6 for a comparison against the DEMNUni void profiles.
@@
Our results may have several applications in upcoming galaxy surveys analysis, in particular concerning DM tracer statistics as the HMF, VSF, and the void-galaxy cross-correlation. As shown in Fig. 3, our multiplicity function provide a much better accuracy than previous modelings when compared with measurements from cosmological simulations. Moreover, it is accurate up and beyond the scale of the lightest clusters (Euclid Collaboration et al., 2019; Hilton et al., 2021) and the smallest voids detectable in galaxy surveys (Contarini et al., 2019). In addition, an accurate modeling of the Lagrangian void density profile can help improving the modeling of redshift space distortions around voids (Pisani et al., 2014; Hamaus et al., 2014, 2016; Cai et al., 2016; Nadathur & Percival, 2019; Nadathur et al., 2019; Hamaus et al., 2020; Nadathur et al., 2020; Hamaus et al., 2022; Woodfinden et al., 2022; Radinović et al., 2023). These results are particular relevant for data analyses from ongoing and forthcoming galaxy redshift surveys, such as Euclid (Laureijs et al., 2011), Roman (Spergel et al., 2015), SPHEREx (Doré et al., 2018), DESI (DESI Collaboration et al., 2016), PFS (Tamura et al., 2016), etc., where the amount of data and the galaxy density will require high precision and accuracy in the theoretical modeling, from the smaller to the larger scales. In this respect, in a future work we plan to further explore the mapping from the Lagrangian to the Eulerian space at scales smaller than $\sim 40h^{-1}{\rm Mpc}$ at $z=0$, on which the spherical mapping already hold within $1\sigma$.
@@
GV acknowledges NASA grant RSA 1692291. GV and AP acknowledge support from the Simons Foundation to the Center for Computational Astrophysics at the Flatiron Institute. The DEMNUni simulations were carried out in the framework of “The Dark Energy and Massive-Neutrino Universe” project, using the Tier-0 IBM BG/Q Fermi machine and the Tier-0 Intel OmniPath Cluster Marconi-A1 of the Centro Interuniversitario del Nord-Est per il Calcolo Elettronico (CINECA). We acknowledge a generous CPU and storage allocation by the Italian Super-Computing Resource Allocation (ISCRA) as well as from the coordination of the “Accordo Quadro MoU per lo svolgimento di attività congiunta di ricerca Nuove frontiere in Astrofisica: HPC e Data Exploration di nuova generazione”, together with storage from INFN-CNAF and INAF-IA2.
Appendix A Standard deviation of the derivative of the filtered density contrast field

### v2

@@ 5.1 The cosmic void case
i) Our VSF model with the effective moving barrier, Eq. (4.1), using the exact numerical multiplicity function (solid blue) from Eqs. (3.4)–(3.5) and the moving barrier of Eq. (4.1) with the MCMC best-fit parameters obtained in Lagrangian space, Eq. (4.2). The blue dashed line shows the corresponding analytical approximation, Eq. (3.14), while the dotted blue line represents the approximated integration, Eq. (3.2). They are mapped in Eulerian space assuming a spherical symmetric evolution, Eq. (5.2).
@@
ii) The SVdW model (green curve) with linear formation threshold $\delta_{\rm v}=-0.623$, mapped in Eulerian space via Eq. (5.2).
@@
iii) The Jennings at al. 2013 [13] VSF model (hereafter Vdn) with linear formation threshold $\delta_{\rm v}=-0.623$ (pink curve). This model, also known as the volume-conserving model, is a Eulerian mapping of the SVdW, which relies mainly on one assumption: the total volume fraction of voids is conserved from Lagrangian to Eulerian space, at the expense of the conservation of the number of voids. This assumption was introduced to avoid normalisation problems in the sharp-$k$ multiplicity function [SVdW] mapped in Eulerian space assuming spherical symmetric evolution.
@@
Figure 5: Eulerian HMF. Top panel: HMF measured from DEMNUni at $z=0$ (black dots) and corresponding Poissonian uncertainty (errorbars); Eulerian spherical mapping of the effective barrier approach implemented in this work (blue), Eqs. (3.2), (4.1), fitting the halo threshold from Eq. (4.2); SMT best-fit model (orange); 4-parameters Tinker et al. 2008 [63] best-fit model. Bottom panel: relative difference between the theoretical models of the upper panel (same colour libelling) with respect to HMF measured in the DEMNUni in unit of Poissonian uncertainty. Errorbars and hatched area show ±$\pm 1\sigma$ interval, dashed area shows the ±$\pm 1\%$ uncertainty
@@
Finally, the lower panel of Fig. 4 shows the residuals at $z=0$ of the above Eulerian VSF models with respect to the VSF measured from DEMNUni (dark colours) and HR-DEMNUni (transparent). Errorbars represent the Poissonian uncertainty. It can be observed that the Eulerian mapping of SVdW and the Vdn model is not able to reproduce the VSF from simulations. Concerning our model, it can be noticed that the spherical mapping is accurate within $1\sigma$ at both large and intermediate scales, while it becomes less accurate at lower scales, on which the spherical Eulerian map over-predicts the number of voids by about $\sim 50\%$. The accuracy at larger scales is in agreement with the fact that the anisotropic part of the strain of the displacement field around density extremants decreases as the scale increases [2, 64]. In particular, the expected value of the ellipticity is proportional to $\sigma(R)/|\delta_{\rm v}|$ [SMT], which makes the spherical approximation accurate on large scales. On smaller scales, other effects, such as non-sphericity and non-locality, may become relevant. From the same arguments, it follows that the accuracy of the spherical Eulerian mapping increases with the redshift.
5.2 The DM halo case
@@
Lagrangian haloes undergo a full collapse in Eulerian space, which makes the mapping conceptually different from the void case. On the one hand, at each redshift it corresponds a unique linear collapse threshold (see Appendix F for details); on the other hand, the spherical symmetric evolution can hardly be accurate in the strongly nonlinear regime of the collapse. In the literature it is shown that, at the first-order approximation, the non-spherical collapse is reflected in a scale-dependent [SMT] and fuzzy formation threshold [18, 29] in Lagrangian space. In addition, non-linear evolution can break the top-hat Lagrangian to Eulerian mass conservation, $M_{\rm E}(R)=M_{\rm TH}(R)$, as the fluid elements of a Lagrangian spherical overdensity may not all and exclusively collapse in the corresponding Eulerian halo [29]. Finally, above all, there is the problem of estimating which regions in the Universe (or in simulations) are collapsed and eventually virialized, i.e. are DM haloes. Different prescriptions result in different halo finders [65]. This makes an a priori modelling of haloes statistics, such as the HMF, extremely challenging. As a consequence, the halo multiplicity function needs to be fitted according to the different halo finders considered [^4] Note that analogous considerations are valid for void finding algorithms for which the corresponding void definition cannot be easily mapped to the excursion-peak theory one. In that case, exactly as for haloes, the effective barrier describing the statistics needs to be fitted against simulations. [63SMT], and for each different halo finding algorithm, a different effective barrier would correspond to it, representing the statistics of each different halo population. In this perspective, we use the HMF measured at $z=0$ from the DEMNUni simulations to find the corresponding halo formation threshold for Lagrangian haloes, as defined in Sec. 2. In particular, to test the stability of our approach against the used halo-finder, we consider different halo catalogues, such as: Friend-of-Friends (FoF) haloes; subhaloes obtained by post-processing FoF haloes with the Gadget-3 Subfind algorithm (which identifies locally overdense regions, i.e. areas enclosed by an isodensity contour that traverses a saddle point) [66, 67]; and Spherical-Overdensity (SO) halo catalogues corresponding to an overdensity threshold equal to the virial value and obtained by exploiting the SO package within the Subfind algorithm. We found that the barrier slightly changes to accommodate the HMFs of the different catalogues, while variations on the formation threshold values are minor.
Then, according to the spherical mapping approximation, we assume the linear halo formation threshold, $\delta_{\rm h}$, to be scale independent and the Eulerian mass to be the same as the Lagrangian one, $M_{\rm E}(R)=M_{\rm TH}(R)$. Following SMT, we consider the subhalo catalogues and, as explained in Appendix F, we use Eq. (4.2) as a prior to fit the threshold, after substituting $\delta_{\rm v}$ with $\delta_{\rm h}$. We find a redshift independent best-fit $\delta_{\rm h}=2.424\pm^{0.076}_{0.071}$. Moreover the corresponding effective barrier parameters in Eq. (4.1) are found to be redshift independent as well, satisfying the regime in which the HMF can be considered self-similar, confirming its redshift universality (see Appendix F for details). Note that, even if this value seems to be larger than the usual spherical collapse threshold, $\delta_{\rm c}\simeq 1.686$, actually our results confirm previous finding in the literature [18]. In fact, the barrier height corresponding to $\delta_{\rm h}=2.424\pm^{0.076}_{0.071}$ is $\alpha=1.209\pm^{0.008}_{0.007}$, only a $\sim 15\%$ different from the  SMT moving barrier, $\sqrt{0.7}\delta_{\rm c}\sim 1.4$.
@@
Fig. 5 shows the results. The upper panel shows: i) the sub-HMF measured at $z=0$ from DEMNUni (black dots), with the corresponding Poissonian uncertainty (error bars); ii) threshold fit $\delta_{\rm h}$ of theoretical HMF from the moving barrier approach using Eqs. (3.2)–(4.1)–(4.2) (blue curve); iii) the SMT best-fit model (orange curve); iv) 4-parameters Tinker et al. 2008 [63] best-fit model. The bottom panel shows the relative differences between the measurements from simulations and the best-fit models of the upper panel in units of the Poissonian uncertainty. The errorbars and the hatched area show the ±$\pm 1\sigma$ interval, the dashed area shows the ±$\pm 1\%$ uncertainty. It can be noticed that, on the scales considered, the model presented in this work, Eqs. (3.2)–(4.1)–(4.2), plus the fitting of the linear formation threshold $\delta_{\rm h}$, allow us to reach a better accuracy than for the SMT best-fit model and similar to the 4-parameters Tinker et al. 2008 [63] best-fit model. Appendix F shows the redshift dependence of all these theoretical HMFs.
@@ 6 The void density profile from the effective barrier
In this Section, we explore the effective barrier approach for statistics beyond the multiplicity function, by considering the void density profile of Lagrangian voids. Note that this is the integral of the void-matter cross-correlation function. Here we focus on the density profile of voids because, beyond the motivations explained at the beginning of Sec. 4, halo collapse destroys the one-to-one map between the Lagrangian and Eulerian profile, which is not the case for voids. Moreover, a theoretical modelling of the Lagrangian density profile of voids would be useful for their analysis in the redshift space [68, 69, 70, 71, 72].
@@
In the standard excursion-set approach, each stochastic realisation of the $\delta({\bf q},R)$ field, given by Eqs. (3.1), statistically corresponds to a realisation of the integrated density profile with respect to a random Lagrangian position ${\bf q}$. Here, we apply this feature using our effective barrier approach to void formation, in order to verify whether it can reproduce the density profile around minima. As the effective barrier $B$ in Eqs. (4.1)–(4.2) does not correspond to the void formation threshold $\delta_{\rm v}$, the corresponding density profile realisations would not describe the physical ones. We avoid this effect by considering two steps. First, at scales larger than the crossing radius, $R_{\times}$, we normalise the field to the actual void formation threshold as
$\delta(R)\rightarrow\delta(R)\,\delta_{\rm v}/B(R)\quad{\rm for}\quad R\geq R_{\times}.$ (6.1)
@@
This step aims to rescale the amplitude of the field from the effective to the physical one. Second, at scales smaller than the crossing radius, we shift the stochastic realisation of the field as
$\delta(R)\rightarrow\delta(R)-B(R_{\times})+\delta_{\rm v}\quad{\rm for}\quad R<R_{\times}.$ (6.2)
@@
This shifts the stochastic realisation $\delta(R)$ by an offset chosen to match the physical void formation threshold at the crossing radius, $\delta(R_{\times})=\delta_{\rm v}$. After crossing, the field behaves as a random walk pinned at the coordinate $(R_{\times},\delta_{\rm v})$ in the $R-\delta$ plane and therefore satisfies the above equation. It follows that the width of the distribution of void profiles is approximated by $\sqrt{s-S(R_{\times})}$ as $s$ increases ($R$ decreases).
@@
Figure 6: Lagrangian density profile, linearly evolved at $z=0$, for voids with linear formation threshold of $\delta_{\rm v}=-0.388$ and void radius between $15-21$Mpc (purple), $33-39$Mpc (red), $58-65$Mpc (green), and $(72-79)$ $h^{-1}$pc (blue). Solid lines show the excursion-set prediction with the effective moving barrier obtained as described in 6; light hatched areas the shortest interval subtending 68% of the distribution; dashed lines show the mean void density profile measured in simulations ICs; light shaded areas the shortest interval subtending 68% of the distribution.
@@
Fig. 6 shows the comparison between the void density profile measured in simulations ICs and linearly evolved at $z=0$ (dashed lines), with respect to stochastic realisations of $\delta(R)$, obtained with the corresponding effective moving barrier, Eqs. (4.1)–(4.2), rescaled via Eqs. (6.1)–(6.2) (solid lines). Different colours represent voids with different formation radii, as specified in the caption. The shaded areas show the shortest interval subtending $68\%$ of the distribution of density profiles in simulations ICs; while the hatched areas show the same quantity for stochastic realisations. Interestingly, in both cases, the distributions are well described by a Gaussian PDF for each radius $R$. The correspondence between simulations ICs and stochastic realisations is very good, in particular: i) the void profile distribution obtained via the effective barrier approach closely matches the one from simulations ICs, as can be noted by the shaded and hatched areas in Fig. 6; ii) the mean of the excursion-set profile distribution reproduces the global behaviour of the mean profile measured in the simulations ICs, while the accuracy depends on the void scale considered; iii) the change of slope around the crossing radius is accurately reproduced; iv) interestingly, the effective barrier predicts the formation of a non-collapsing overdensity in which smaller voids are embedded, as observed in simulations ICs.
@@
Our void profile prediction in Lagrangian space is the third main result of this work, and the methodology to model it has been presented here for the first time in the literature. In particular, this modelling uses the same effective barrier of Eqs. (4.1)–(4.2) used to derive the theoretical multiplicity function and VSF presented in this work. This modelling would not be representative of simulation ICs measurements unless the effective barrier contained statistical information beyond the multiplicity function. In this work we do not explore the Lagrangian to Eulerian mapping for the void density profile, which we aim to study in a future work. However, it has already been shown in the literature that it is possible to obtain the Eulerian void density profile from the Lagrangian one analytically  [30, 73] and we think that similar techniques can be applied also to our formalism.
@@ 7 Conclusions
In this work, we show that using an effective moving barrier in the excursion-set framework, it is possible to combine the latter with the peak theory in Lagrangian space. In particular: i) we provide in Sec. 2 a precise halo and void formation definition in Lagrangian space, that was missing for voids, which combines excursion-set and peak theory; ii) in Sec. 3, for a general moving barrier in the presence of scale correlations, we derive Eqs. (3.14) and (3.2), i.e. a multiplicity function able to recover both halo and void statistics from small to very large scales $S$. In particular, we show that the exact computation of $\langle\delta^{\prime 2}_{S}\rangle$ is crucial to obtain accurate analytical approximations of the general solution (see Fig. 1 for a detailed comparison with the numerical multiplicity function obtained via MC realisations); iii) in Sec. 4 we derive, in the void case, the parameters characterising the SMT-like effective moving barrier, in Eq. (4.1), as a function of the void formation threshold, $\delta_{\rm v}$, by inserting Eq. (4.1) in Eq. (3.5), Eq. (3.14) and Eq. (3.2), and fitting the VSF obtained via the latter against the Lagrangian VSF measured in the DEMNUni ICs; iv) in Sec. 5 we discuss the Lagrangian to Eulerian mapping, in particular we implement the spherical map in Eulerian space for the Lagrangian VSF and HMF, introduced in Secs. 34. We obtain a good agreement with simulations for intermediate and large voids and we accurately reproduce the HMF by fitting the linear halo formation threshold from simulations; v) in Sec. 6 we show that, in the excursion-set framework with the effective barrier, it is possible to derive the Lagrangian void density profile, using the effective barrier of Eqs. (4.1)–(4.2) entering the corresponding multiplicity function, Eq. (3.5), derived in this work. See Fig. 6 for a comparison against the DEMNUni void profiles.
@@
Our results may have several applications in upcoming galaxy surveys analysis, in particular concerning DM tracer statistics as the HMF, VSF, and the void-galaxy cross-correlation. As shown in Fig. 3, our multiplicity function provides a much better accuracy than previous modellings when compared with measurements from cosmological simulations. Moreover, it is accurate up and beyond the scale of the lightest clusters [74, 75] and the smallest voids detectable in galaxy surveys [76]. In addition, an accurate modelling of the Lagrangian void density profile can help improving the modelling of redshift space distortions around voids [68, 77, 70, 71, 78, 79, 80, 81, 82, 83, 72]. These results are particularly relevant for data analyses from ongoing and forthcoming galaxy redshift surveys, such as Euclid [84], Roman [85], SPHEREx [86], DESI [87], PFS [88], etc., where the amount of data and the galaxy density will require high precision and accuracy in the theoretical modelling, from the smaller to the larger scales. In this respect, note that in this work we considered around $2\times 10^{5}$ voids (see Tab 2), while, for example, in Euclid the number of expected voids is around 6000 [28]. This makes the modelling presented in this work very accurate, with much smaller uncertainties than those expected from future surveys. In this respect, in a future work we plan to further explore the mapping from the Lagrangian to the Eulerian space at scales smaller than $\sim 40h^{-1}{\rm Mpc}$ at $z=0$, on which the spherical mapping already holds within $1\sigma$. For voids in galaxy surveys other observational effects beyond the Lagrangian to Eulerian mapping in real space should be considered, such as the redshift-space distortion impact. We plan to explore them in future works by applying well known techniques in the literature [e.g. 89, 90, 71, 78, 91, 92].
@@
GV would like to thank Nickolas Kokron for useful discussions. GV acknowledges NASA grant EUCLID12-0004. GV and AP acknowledge support from the Simons Foundation to the Center for Computational Astrophysics at the Flatiron Institute. The DEMNUni simulations were carried out in the framework of “The Dark Energy and Massive-Neutrino Universe” project, using the Tier-0 IBM BG/Q Fermi machine and the Tier-0 Intel OmniPath Cluster Marconi-A1 of the Centro Interuniversitario del Nord-Est per il Calcolo Elettronico (CINECA). We acknowledge a generous CPU and storage allocation by the Italian Super-Computing Resource Allocation (ISCRA) as well as from the coordination of the “Accordo Quadro MoU per lo svolgimento di attività congiunta di ricerca Nuove frontiere in Astrofisica: HPC e Data Exploration di nuova generazione”, together with storage from INFN-CNAF and INAF-IA2.
Appendix A Correlation of the density contrast field scale derivatives